# Classifying localizing subcategories of a Grothendieck category

### v1

@@ Abstract.
Let $\mathcal{A}$ be a Grothendieck category, with $\operatorname{Sp}\mathcal{A}$ denoting the set of isomorphism classes of indecomposable injective objects of $A$, and $\operatorname{ASpec}\mathcal{A}$ the atom spectrum of $\mathcal{A}$. In this paper, we introduce new topologies on $\operatorname{Sp}\mathcal{A}$ and $\operatorname{ASpec}\mathcal{A}$, and classify localizing subcategories of:
(i) locally monoform categories,
(ii) locally finitely generated categories, and
(iii) locally coherent categories
via open subsets of $\operatorname{Sp}\mathcal{A}$ and $\operatorname{ASpec}\mathcal{A}$ under respective topologies. As an application, we specialize our investigation to the case of commutative (coherent) rings.
@@
Atom spectrum; finite type; localizing subcategory; Serre subcategoy; Ziegler spectrum.
@@ 1. Introduction
Throughout this paper, we assume that $\mathcal{A}$ is a Grothendieck category. The set $\operatorname{Sp}\mathcal{A}$ consisting of isomorphism classes of indecomposable injective objects of $A$, plays a central role in classifying localizing subcategories of a Grothendieck category (see [G, Kr]). Over a commutative noetherian ring $\mathcal{A}$, there is a bijective correspondence between $\operatorname{Spec}A$ and $\operatorname{Sp}\mathcal{A}$.
@@
Over a locally coherent category $\mathcal{A}$, Herzog [H] defined the Ziegler spectrum of $\mathcal{A}$, denoted $\operatorname{Zg}\mathcal{A}$, as an alternative topology on $\operatorname{Sp}\mathcal{A}$. This topology was originally defined by Ziegler [Z] associating to a ring $A$, a topological space whose points are the isomorphism classes of the indecomposable pure-injective $A$-modules.
[…]
In this paper, we define specific topologies on $\operatorname{Sp}\mathcal{A}$ and $\operatorname{ASpec}\mathcal{A}$ for certain Grothendieck categories $\mathcal{A}$. Subsequently, we classify localizing subcategories of categories $\mathcal{A}$ using open subsets of these topologies.
@@
In Section 2, we investigate categories with enough atoms. An abelian category $\mathcal{A}$ is said to be locally monoform if every nonzero object in $\mathcal{A}$ contains a monoform subobject. Unfortunately, Grothendieck categories are not generally locally monoform, which limits our ability to find out further insights about $\mathcal{A}$ via $\operatorname{ASpec}\mathcal{A}$. A well-known example of locally monoform categories are locally noetherian categories ; see [K]. For any subcategory $\mathcal{X}$ of $\mathcal{A}$ and any subset $\mathcal{U}$ of $\operatorname{ASpec}\mathcal{A}$, we define $\operatorname{ASupp}\mathcal{X}=\bigcup_{M\in\mathcal{X}}\operatorname{ASupp}M$ and $\operatorname{ASupp}^{-1}\mathcal{U}=\{M\in\mathcal{A}|\hskip 2.84544pt\operatorname{ASupp}M\subseteq\mathcal{U}\}$. The following theorem classifies localizing subcategories of a locally monoform category:
[…]
In this section, we also prove that semi-noetherian categories are locally monoform; see section 2.
A Grothendieck category $\mathcal{A}$ is said to be locally finitely generated if $\mathcal{A}$ has a generating set of finitely generated objects. The following theorem is one of the main theorem of Section 3:
[…]
Let $\mathcal{A}$ be a locally finitely generated category. Then the map $\mathcal{X}\mapsto\operatorname{ASupp}\mathcal{X}$ provides a bijective correspondence between the lattice of localizing subcategories of $\mathcal{A}$ and the lattice of open subsets of $\operatorname{ASpec}\mathcal{A}$. The inverse map is given by $\mathcal{U}\mapsto\operatorname{ASupp}^{-1}\mathcal{U}$.
For an object $M$ in $\mathcal{A}$, we define the Ziegler support of $M$ that is $\operatorname{ZSupp}M=\{I\in\operatorname{Zg}\mathcal{A}|\hskip 2.84544pt\operatorname{Hom}(M,I)\neq 0\}$. We show that $\{\operatorname{ASupp}M|\hskip 2.84544ptM\in\mathcal{A}\}$ forms a basis of open subsets for a topology on $\operatorname{Sp}\mathcal{A}$. For any subcategory $\mathcal{X}$ of $\mathcal{A}$, we define $\operatorname{ZSupp}\mathcal{X}=\bigcup_{M\in\mathcal{X}}\operatorname{ZSupp}M$ and for any subset $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$, we define $\operatorname{ZSupp}^{-1}\mathcal{U}=\{M\in\mathcal{A}|\hskip 2.84544pt\operatorname{ZSupp}M\subset\mathcal{U}\}$. The following theorem classifies localizing subcategories of a locally finitely generated category $\mathcal{A}$ using open subsets of $\operatorname{Sp}\mathcal{A}$.
[…]
Let $\mathcal{A}$ be a locally finitely generated category. The map $\mathcal{X}\mapsto\operatorname{ZSupp}\mathcal{X}$ provides a bijective correspondence between the lattice of localizing subcategories of $\mathcal{A}$ and the lattice of open subsets of $\operatorname{Sp}\mathcal{A}$. The inverse map is given by $\mathcal{U}\mapsto\operatorname{ZSupp}^{-1}\mathcal{U}$.
@@
A Grothendieck category $\mathcal{A}$ is said to be locally coherent if $\mathcal{A}$ has a generating set of finitely presented objects and the full subcategory $\operatorname{fp-}\mathcal{A}$ of finitely presented objects in $\mathcal{A}$ is abelian. In Section 4, we assume that $\mathcal{A}$ is locally coherent. We define a new topology on $\operatorname{ASpec}\mathcal{A}$ in which $\{\operatorname{ASupp}M|\hskip 2.84544ptM\in\operatorname{fp-}\mathcal{A}\}$ forms a basis of open subsets of $\operatorname{ASpec}\mathcal{A}$. We use the symbol $\operatorname{ZASpec}\mathcal{A}$ instead of $\operatorname{ASpec}\mathcal{A}$ with this topology. The following theorem classifies localizing subcategories of a locally coherent category:
[…]
@@
As a conclusion, we prove that there is a bijective correspondence between the lattice of open subsets of $\operatorname{ZASpec}\mathcal{A}$ and the lattice of localizing subcategories of finite type of $\mathcal{A}$; see section 4. Furthermore, we show that a localizing subcategory $\mathcal{X}$ is of finite type of $\mathcal{A}$ if and only if $\operatorname{ASupp}\mathcal{X}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$. We characterize quasi-compact open subsets of $\operatorname{ZASpec}\mathcal{A}$; see section 4. As $\operatorname{fp-}\mathcal{A}$ is abelian, $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ can be investigated independently. We denote the atoms in $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ by $\alpha_{f}=\overline{H}$, where $H$ is a monoform object in $\operatorname{fp-}\mathcal{A}$. It is shown that $H$ is also a monoform object in $\mathcal{A}$ so that $\alpha=\overline{H}$ is an atom in $\operatorname{ASpec}\mathcal{A}$. Then the map $\alpha_{f}\stackrel{{\scriptstyle g}}{{\mapsto}}\alpha$ embeds $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ as a subspace of $\operatorname{ASpec}\mathcal{A}$. We prove that if $\mathcal{A}$ is locally monoform then $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ is dense in $\operatorname{ASpec}\mathcal{A}$; see section 4. It is known as Cohen’s theorem that if prime ideals of a commutative ring $A$ are finitely generated, then $A$ is noetherian. In the following theorem we extend this theorem for locally coherent categories.
[…]
@@
If $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}=\operatorname{ASpec}\mathcal{A}$, then $\mathcal{A}$ is locally noetherian.
@@
For any object $M$ in $\mathcal{A}$, we define $\operatorname{fASupp}M=\{\alpha_{f}=\overline{H}\in\operatorname{ASpec}\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\alpha=\overline{H}\in\operatorname{ASupp}M\}.$ For any subcategory $\mathcal{X}$ of $\mathcal{A}$ , we define $\operatorname{fASupp}\mathcal{X}=\bigcup_{M\in\mathcal{X}}\operatorname{fASupp}M$. In the following theorem, we provide a sufficient condition under which a localizing subcategory of $\mathcal{A}$ is of finite type.
[…]
In Section 5, we investigate Ziegler spectrum of a locally coherent category. Herzog showed that the set $\{\operatorname{ZSupp}M|\hskip 2.84544ptM\in\operatorname{fp-}\mathcal{A}\}$ can serve as a basis for a topology on $\mathcal{A}$. This topology, denoted by $\operatorname{Zg}\mathcal{A}$, is known as Ziegler topology. The following theorem provides a classification of localizing subcategories of finite type of $\mathcal{A}$ via open subsets of $\operatorname{Zg}\mathcal{A}$:
[…]
Let $\mathcal{A}$ be a locally coherent category. Then the map $\mathcal{X}\mapsto\operatorname{ZSupp}\mathcal{X}$ provides a bijective correspondence between the lattice of localizing subcategories of finite type of $\mathcal{A}$ and the lattice of open subsets of $\operatorname{Zg}\mathcal{A}$. The inverse map is given by $\mathcal{U}\mapsto\operatorname{ZSupp}^{-1}\mathcal{U}$.
As a conclusion of the above theorem, a localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ is of finite type if and only if $\operatorname{ZSupp}\mathcal{X}$ is an open subset of $\operatorname{Zg}\mathcal{A}$.
In Section 6, we assume that $A$ is a commutative ring. We denote $\operatorname{ASpec}\operatorname{Mod}A$ by $\operatorname{ASpec}A$ and we show that =$\mathcal{I}=\{\operatorname{ASupp}A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}$ is an ideal of $\mathcal{A}\}$ forms a basis for $\operatorname{ASpec}A$. We establish a bijective correspondence between the lattices of specialization-closed subsets of $\operatorname{Spec}A$ and the lattice of the open subsets of $\operatorname{ASpec}A$. We denote $\operatorname{Sp}\operatorname{Mod}A$ by $\operatorname{Sp}A$. There are many classifications of Gabriel topologies on $A$ (e.g., see [G, St]). In the following theorem, we provide an alternative classification:
[…]
The map $\mathfrak{F}\stackrel{{\scriptstyle D}}{{\mapsto}}D(\mathfrak{F})$ together with the inverse map $D\stackrel{{\scriptstyle\mathfrak{F}}}{{\mapsto}}\mathfrak{F}_{D}$ establishes a bijective correspondence between the lattices of Gabriel topologies on $A$ and the lattice of closed subsets of $\operatorname{Sp}A$.
@@
Assume that $A$ is a commutative coherent ring. We show that =$\mathcal{F}\mathcal{I}=\{\operatorname{ASupp}A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}$ is a finitely generated ideal of $A\}$ forms a basis for $\operatorname{ZASpec}A$. For a spectral topological space $X$, Hochster [Ho] endows the underlying set with a new, dual topological by defining its open subsets as those of the form $Y=\bigcup_{i\in\varOmega}Y_{i}$, where $X\setminus Y_{i}$ is a quasi-compact open subset of $X$ for each $i\in\varOmega$. The symbol $X^{\star}$ denotes $X$ with the new topology. We write $\operatorname{Spec}^{\star}A$ for $(\operatorname{Spec}A)^{\star}$. We establish a bijective correspondence between the lattices of open subsets of $\operatorname{ZASpec}^{\star}A$ and those of $\operatorname{ZASpec}A$.
@@ 2. Categories with enough atoms
[…]
@@
[…]
@@
In view of [Sto, p.631], for a commutative ring $A$, there is a bijection between $\operatorname{ASpec}A$ and $\operatorname{Spec}A$. A subset $\varPhi$ of $\operatorname{Spec}A$ is said to be specialization-closed if for any prime ideals $\mathfrak{p}$ and $\mathfrak{q}$ of $A$ with $\mathfrak{p}\subseteq\mathfrak{q}$, the condition $\mathfrak{p}\in\varPhi$ implies that $\mathfrak{q}\in\varPhi$. A corresponding subset in $\operatorname{ASpec}\mathcal{A}$ can be defined as follows
[…]
@@
[…]
@@
[…]
For a Serre subcategory $\mathcal{X}$ of $\mathcal{A}$, we define the quotient category $\mathcal{A}/\mathcal{X}$ in which the objects are those in $\mathcal{A}$ and for objects $M$ and $N$ in $\mathcal{A}$, we have
$\operatorname{Hom}_{\mathcal{A}/\mathcal{X}}(M,N)=\underset{(M^{\prime},N^{\prime})\in\mathcal{S}_{M,N}}{\underrightarrow{\rm lim}}\operatorname{Hom}_{\mathcal{A}}(M^{\prime},N/N^{\prime})$
where $\mathcal{S}_{M,N}$ is a directed set defined by
$\mathcal{S}_{M,N}=\{(M^{\prime},N^{\prime})|M^{\prime}\subset M,N^{\prime}\subset N\hskip 2.84544pt{\rm with}\hskip 3.69899ptM/M^{\prime},N^{\prime}\in\mathcal{X}\}.$
We observe that $\mathcal{A}/\mathcal{X}$ is a Grothendieck category together with a canonical exact functor $F:\mathcal{A}\rightarrow\mathcal{A}/\mathcal{X}$. We refer the reader to [G] or [Po, Chap 4] for more details and the basic properties of the quotient categories.
@@
The following lemma is crucial to prove the main theorem of this section.
[…]
@@
We present the main result of this section.
[…]
@@
For a localizing subcategoriy $\mathcal{X}$ of $\mathcal{A}$ and an open subset $\mathcal{U}$ of $\operatorname{ASpec}\mathcal{A}$, it suffices to prove that $\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})=\mathcal{X}$ and $\operatorname{ASupp}(\operatorname{ASupp}^{-1}\mathcal{U})=\mathcal{U}$. But this follows easily from the definition and using section 2. ∎
@@
Let us assume that $\sigma=\rho+1$ and denote by $F_{\rho}:\mathcal{A}\rightarrow\mathcal{A}/\mathcal{A}_{\rho}$ the canonical functor and by $G_{\rho}:\mathcal{A}/\mathcal{A}_{\rho}\rightarrow\mathcal{A}$ the right adjoint functor of $F_{\rho}$. Then an object $X$ in $\mathcal{A}$ will belong to $\mathcal{A}_{\sigma}$ if and only if $F_{\rho}(X)\in{\rm Ob}(\mathcal{A}/\mathcal{A}_{\rho})_{0}$. The left exact radical functor (torsion functor) corresponding to $\mathcal{A}_{\rho}$ is denoted by $t_{\rho}$. If $\sigma$ is a limit ordinal, then $\mathcal{A}_{\sigma}$ is the localizing subcategory generated by all localizing subcategories $\mathcal{A}_{\rho}$ with $\rho<\sigma$. It is clear that if $\sigma\leq\sigma^{\prime}$, then $\mathcal{A}_{\sigma}\subseteq\mathcal{A}_{\sigma^{\prime}}$. Moreover, since the class of all localizing subcategories of $\mathcal{A}$ is a set, there exists an ordinal $\tau$ such that $\mathcal{A}_{\sigma}=\mathcal{A}_{\tau}$ for all $\sigma\leq\tau$. Let us put $\mathcal{A}_{\tau}=\cup_{\sigma}\mathcal{A}_{\sigma}$. Then $\mathcal{A}$ is said to be semi-noetherian if $\mathcal{A}=\mathcal{A}_{\tau}$. We also say that the localizing subcategories $\{\mathcal{A}_{\sigma}\}_{\sigma}$ define the Krull-Gabriel filtration of $\mathcal{A}$. We say that an object $M$ in $\mathcal{A}$ has the Krull-Gabriel dimension defined or $M$ is semi-noetherian if $M\in{\rm Ob}(\mathcal{A}_{\tau})$. The smallest ordinal $\sigma$ so that $M\in{\rm Ob}(\mathcal{A}_{\sigma})$ is denoted by $\operatorname{KG-dim}M$. Because the class of ordinals is well-ordered, throughout this paper, $\omega$ is denoted the smallest limit ordinal. We observe that $\operatorname{KG-dim}0=-1$ and $\operatorname{KG-dim}M\leq 0$ if and only if $\operatorname{ASupp}M\subseteq\textnormal{m-}\operatorname{ASpec}\mathcal{A}$.
@@
[…]
Given an ordinal $\sigma\geq 0$, we recall from [MR or GW] that an object $M$ in $\mathcal{A}$ is $\sigma$-critical provided $\operatorname{KG-dim}M=\sigma$ while $\operatorname{KG-dim}M/N<\sigma$ for all nonzero subobjects $N$ of $M$. It is clear that any nonzero subobject of a $\sigma$-critical object is $\sigma$-critical. An object $M$ is said to be critical if it is $\sigma$-critical for some ordinal $\sigma$. It is clear to see that any critical object is monoform.
[…]
@@
The following lemma is crucial in this section.
[…]

### v2

@@ Abstract.
Let $\mathcal{A}$ be a locally coherent Grothendieck category, $\operatorname{fp-}\mathcal{A}$ be the full subcategory of $\mathcal{A}$ consisting of finitely presented objects and $\operatorname{ASpec}\mathcal{A}$ be the atom spectrum of $\mathcal{A}$. In this paper, we classify localizing subcategories of finite type of $\mathcal{A}$ via open subsets of $\operatorname{ASpec}\mathcal{A}$. We investigate $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ and show that if $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}=\operatorname{ASpec}\mathcal{A}$, then $\mathcal{A}$ is locally noetherian. As an application, we specialize our investigation to the case of commutative coherent rings.
@@
Atom spectrum; coherent ring, localizing subcategory; Serre subcategoy; Ziegler spectrum.
@@ 1. Introduction
Throughout this paper, we assume that $\mathcal{A}$ is a Grothendieck category. The idea of classifying subcategories of $\mathcal{A}$ using an appropriate spectrum originates from the classification of subcategories of modules over a commutative ring $A$ using $\operatorname{Spec}A$, which has been established by numerous authors (e.g. [Hop, Hov, GP, T]). The initial work in extending this concept to Grothendieck categories was carried out by Gabriel [G], who defined the spectrum $\operatorname{Sp}\mathcal{A}$ as the set of isomorphism classes of indecomposable injective objects in $\mathcal{A}$. In the case of a commutative noetherian ring, these indecomposable injective modules are in one-to-one correspondence with prime ideals.
@@
For a locally coherent category $\mathcal{A}$, Herzog [H] defined $\operatorname{Zg}\mathcal{A}$, the Ziegler spectrum of $\mathcal{A}$, which imposes an alternative topology on $\operatorname{Sp}\mathcal{A}$. This topology was originally defined by Ziegler [Z] associating to a ring $A$, a topological space whose points are the isomorphism classes of indecomposable pure-injective $A$-modules.
Recently, Kanda [K] introduced the atom spectrum $\operatorname{ASpec}\mathcal{A}$, for an abelian category $\mathcal{A}$ that may not have enough injective objects, and endowed it with a topology. This construction was inspired by Storrer’s [St] work on monoform modules and their equivalence relation over non-commutative rings. For an object $M$ in $\mathcal{A}$, the atom support $\operatorname{ASupp}M$ can be defined analogously to the support of a module over a commutative ring. This paper is devoted to classifying the localizing subcategories of finite type of a locally coherent category $\mathcal{A}$ via the open subsets of $\operatorname{ASpec}\mathcal{A}$.
@@
Unfortunately, Grothendieck categories does not have generally enough atoms, which limits our ability to find out further insights about $\mathcal{A}$ via $\operatorname{ASpec}\mathcal{A}$. A well-known example of a categories with enough atoms is locally noetherian categories; see [K]. An abelian category $\mathcal{A}$ is said to be locally monoform if every nonzero object in $\mathcal{A}$ contains a monoform subobject. In Section 2, we show that if $\mathcal{A}$ is a locally monoform category, then there is a bijective correspondence between the lattice of localizing subcategories of $\mathcal{A}$ and the lattice of open subsets of $\operatorname{ASpec}\mathcal{A}$; see section˜2 . Furthermore, we also show that semi-noetherian categories are locally monoform; see section˜2.
@@
A Grothendieck category $\mathcal{A}$ is said to be locally coherent if $\mathcal{A}$ has a generating set of finitely presented objects and the full subcategory $\operatorname{fp-}\mathcal{A}$ of finitely presented objects in $\mathcal{A}$ is abelian. In Section 3, we assume that $\mathcal{A}$ is locally coherent. We define a new topology on $\operatorname{ASpec}\mathcal{A}$ in which $\{\operatorname{ASupp}M|\hskip 2.84544ptM\in\operatorname{fp-}\mathcal{A}\}$ forms a basis of open subsets of $\operatorname{ASpec}\mathcal{A}$. We use the symbol $\operatorname{ZASpec}\mathcal{A}$ instead of $\operatorname{ASpec}\mathcal{A}$ with this topology. For any subcategory $\mathcal{X}$ of $\mathcal{A}$ and any subset $\mathcal{U}$ of $\operatorname{ASpec}\mathcal{A}$, assume that $\operatorname{ASupp}\mathcal{X}=\bigcup_{M\in\mathcal{X}}\operatorname{ASupp}M$ and $\operatorname{ASupp}^{-1}\mathcal{U}=\{M\in\mathcal{A}|\hskip 2.84544pt\operatorname{ASupp}M\subseteq\mathcal{U}\}$. The following theorem classifies Serre subcategories of $\operatorname{fp-}\mathcal{A}$ using open subsets of $\operatorname{ZASpec}\mathcal{A}$:
Theorem 1.1 (section˜3).
@@
As a conclusion, we prove that there is a bijective correspondence between the lattice of open subsets of $\operatorname{ZASpec}\mathcal{A}$ and the lattice of localizing subcategories of finite type of $\mathcal{A}$; see section˜3. Furthermore, we show that a localizing subcategory $\mathcal{X}$ is of finite type of $\mathcal{A}$ if and only if $\operatorname{ASupp}\mathcal{X}$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$. As $\operatorname{fp-}\mathcal{A}$ is abelian, $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ can be investigated independently. It is known as Cohen’s theorem that if prime ideals of a commutative ring $A$ are finitely generated, then $A$ is noetherian. In the following theorem we extend this theorem for locally coherent categories:
Theorem 1.2 (section˜3).
@@
Let $\mathcal{A}$ be a locally coherent category. If $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}=\operatorname{ASpec}\mathcal{A}$, then $\mathcal{A}$ is locally noetherian.
@@
We denote the atoms in $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ by $\alpha_{f}=\overline{H}$, where $H$ is a monoform object in $\operatorname{fp-}\mathcal{A}$. It is shown that $H$ is also a monoform object in $\mathcal{A}$ so that $\alpha=\overline{H}$ is an atom in $\operatorname{ASpec}\mathcal{A}$. Then the map $\alpha_{f}\stackrel{{\scriptstyle g}}{{\mapsto}}\alpha$ embeds $\operatorname{ASpec}\operatorname{fp-}\mathcal{A}$ as a subspace of $\operatorname{ASpec}\mathcal{A}$. Then for any object $M$ in $\mathcal{A}$, we define $\operatorname{fASupp}M=\{\alpha_{f}=\overline{H}\in\operatorname{ASpec}\operatorname{fp-}\mathcal{A}|\hskip 2.84544pt\alpha=\overline{H}\in\operatorname{ASupp}M\}.$ For any subcategory $\mathcal{X}$ of $\mathcal{A}$, we define $\operatorname{fASupp}\mathcal{X}=\bigcup_{M\in\mathcal{X}}\operatorname{fASupp}M$. In the following theorem, we provide a sufficient condition under which a localizing subcategory of $\mathcal{A}$ is of finite type:
Theorem 1.3 (section˜3).
Let $\mathcal{A}$ be a locally coherent category. Then any localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ with $\operatorname{fASupp}\mathcal{X}=\operatorname{ASupp}\mathcal{X}$ is of finite type.
For any object $M$ in $\mathcal{A}$, set $\operatorname{\mathcal{O}}(M)=\{I\in\operatorname{Zg}\mathcal{A}|\hskip 2.84544pt\operatorname{Hom}(M,I)\neq 0\}.$ Herzog showed that over a locally coherent category, the set $\{\operatorname{\mathcal{O}}(M)|\hskip 2.84544ptM\in\operatorname{fp-}\mathcal{A}\}$ can serve as a basis for a topology on $\mathcal{A}$. This topology, denoted by $\operatorname{Zg}\mathcal{A}$, is known as Ziegler topology of $\mathcal{A}$. A category $\mathcal{A}$ is said to be locally uniform if every nonzero object in $\mathcal{A}$ has a uniform subobject. In Section 4, we show that $\mathcal{A}$ is locally monoform if and only if $\mathcal{A}$ is locally uniform and the map $\theta:\operatorname{ZASpec}\mathcal{A}\rightarrow\operatorname{Zg}\mathcal{A}$, given by $\alpha\mapsto E(\alpha)$ is a homeomorphism. We provide an example which show that any locally coherent is not necessarily locally monoform; see section˜4. Furthermore, we provide a sufficient condition under which a locally coherent category is semi-noetherian (and so a locally monoform); see section˜4.
@@
In Section 5, we assume that $A$ is a commutative coherent ring. We denote $\operatorname{ZASpec}\operatorname{Mod}A$ and $\operatorname{Zg}\operatorname{Mod}A$ by $\operatorname{ASpec}A$ and $\operatorname{Zg}A$. For a spectral topological space $X$, Hochster [Ho] endows the underlying set with a new, dual topological by defining its open subsets as those of the form $Y=\bigcup_{i\in\varOmega}Y_{i}$, where $X\setminus Y_{i}$ is a quasi-compact open subset of $X$ for each $i\in\varOmega$. The symbol $X^{\star}$ denotes $X$ with the new topology. We write $\operatorname{Spec}^{\star}A$ for $(\operatorname{Spec}A)^{\star}$. We establish a bijective correspondence between the lattice of open subsets of $\operatorname{ZASpec}^{\star}A$ and that of $\operatorname{ZASpec}A$; see section˜5. As our main application, we use Gabriel topologies on $A$ to classify the localizing subcategories of modules of finite type over a commutative coherent ring:
Theorem 1.4 (section˜5).
There are bijective correspondence between:
$\rm(1)$} the lattice of open subsets of $\operatorname{ZASpec}A$,
$\rm(2)$} the lattice of closed subsets of $\operatorname{Zg}A$,
$\rm(3)$} the lattice of Gabriel topologies having bases of finitely generated ideals of $A$,
$\rm(4)$} The lattice of localizing subcategory of finite type of $\operatorname{Mod}A$.
@@ 2. Categories with enough atoms
$\overline{H}=\{G\in\operatorname{ASpec}_{0}\mathcal{A}|\hskip 2.84544ptH\hskip 2.84544pt{\rm and}\hskip 2.84544ptG$ have a common nonzero subobject}.\}.
Definition 2.3.
@@
$\operatorname{ASupp}M=\{\alpha\in\operatorname{ASpec}\mathcal{A}|\hskip 2.84544pt{\rm there\hskip 2.84544ptexists\hskip 2.84544pt}H\in\alpha\hskip 2.84544pt{\rm which\hskip 2.84544ptis\hskip 2.84544pta\hskip 2.84544ptsubquotient\hskip 2.84544ptof\hskip 2.84544pt}M\}.$
@@
In view of [Sto, p. 631], for a commutative ring $A$, there is a bijection between $\operatorname{ASpec}A$ and $\operatorname{Spec}A$. A subset $\varPhi$ of $\operatorname{Spec}A$ is said to be specialization-closed if for any prime ideals $\mathfrak{p}$ and $\mathfrak{q}$ of $A$ with $\mathfrak{p}\subseteq\mathfrak{q}$, the condition $\mathfrak{p}\in\varPhi$ implies that $\mathfrak{q}\in\varPhi$. A corresponding subset in $\operatorname{ASpec}\mathcal{A}$ can be defined as follows:
Definition 2.5.
@@
Definition 2.6.
An atom $\alpha$ in $\operatorname{ASpec}\mathcal{A}$ is said to be maximal if there exists a simple object $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$. The class of all maximal atoms in $\operatorname{ASpec}\mathcal{A}$ is denoted by $\textnormal{m-}\operatorname{ASpec}\mathcal{A}$.
Definition 2.7.
@@
Definition 2.8.
@@
The following lemma is crucial to classify the localizing subcategories of a locally monoform category.
Lemma 2.10.
@@
Proposition 2.11.
@@
For a localizing subcategory $\mathcal{X}$ of $\mathcal{A}$ and an open subset $\mathcal{U}$ of $\operatorname{ASpec}\mathcal{A}$, it suffices to prove that $\operatorname{ASupp}^{-1}(\operatorname{ASupp}\mathcal{X})=\mathcal{X}$ and $\operatorname{ASupp}(\operatorname{ASupp}^{-1}\mathcal{U})=\mathcal{U}$. But this follows easily from the definition and using section˜2. ∎
@@
Let us assume that $\sigma=\rho+1$ and denote by $F_{\rho}:\mathcal{A}\rightarrow\mathcal{A}/\mathcal{A}_{\rho}$ the canonical functor and by $G_{\rho}:\mathcal{A}/\mathcal{A}_{\rho}\rightarrow\mathcal{A}$ the right adjoint functor of $F_{\rho}$. Then an object $X$ in $\mathcal{A}$ will belong to $\mathcal{A}_{\sigma}$ if and only if $F_{\rho}(X)\in{\rm Ob}(\mathcal{A}/\mathcal{A}_{\rho})_{0}$. The left exact radical functor (torsion functor) corresponding to $\mathcal{A}_{\rho}$ is denoted by $t_{\rho}$. If $\sigma$ is a limit ordinal, then $\mathcal{A}_{\sigma}$ is the localizing subcategory generated by all localizing subcategories $\mathcal{A}_{\rho}$ with $\rho<\sigma$. It is clear that if $\sigma\leq\sigma^{\prime}$, then $\mathcal{A}_{\sigma}\subseteq\mathcal{A}_{\sigma^{\prime}}$. Moreover, since the class of all localizing subcategories of $\mathcal{A}$ is a set, there exists an ordinal $\tau$ such that $\mathcal{A}_{\sigma}=\mathcal{A}_{\tau}$ for all $\sigma\geq\tau$. Let us put $\mathcal{A}_{\tau}=\cup_{\sigma}\mathcal{A}_{\sigma}$. Then $\mathcal{A}$ is said to be semi-noetherian if $\mathcal{A}=\mathcal{A}_{\tau}$. We also say that the localizing subcategories $\{\mathcal{A}_{\sigma}\}_{\sigma}$ define the Krull-Gabriel filtration of $\mathcal{A}$. We say that an object $M$ in $\mathcal{A}$ has the Krull-Gabriel dimension defined or $M$ is semi-noetherian if $M\in{\rm Ob}(\mathcal{A}_{\tau})$. The smallest ordinal $\sigma$ so that $M\in{\rm Ob}(\mathcal{A}_{\sigma})$ is denoted by $\operatorname{KG-dim}M$. We observe that $\operatorname{KG-dim}0=-1$ and $\operatorname{KG-dim}M\leq 0$ if and only if $\operatorname{ASupp}M\subseteq\textnormal{m-}\operatorname{ASpec}\mathcal{A}$.
@@
Definition 2.13.
Given an ordinal $\sigma\geq 0$, we recall from [MR, GW] that an object $M$ in $\mathcal{A}$ is $\sigma$-critical provided $\operatorname{KG-dim}M=\sigma$ while $\operatorname{KG-dim}M/N<\sigma$ for all nonzero subobjects $N$ of $M$. It is clear that any nonzero subobject of a $\sigma$-critical object is $\sigma$-critical. An object $M$ is said to be critical if it is $\sigma$-critical for some ordinal $\sigma$. It is clear to see that any critical object is monoform.
Lemma 2.14.
@@
Lemma 2.15.